The temperature of an ideal gas remains constant while the absolute pressure changes from 100 kpa to 800 kpa. If initial volume is 100 liters, what is the final volume in liters?




1
Expert's answer
2022-03-23T13:59:59-0400

In the described system, the amount of the gas does not change, the same as the temperature, so we can consider the following statement:

"\frac{P_1\times V_1}{T}= \frac{P_2\times V_2}{T}"

After simplifying it leads to following equation allowing to determine the volume of the gas after the pressure change:

"P_1 \times V_1=P_2\times V_2"

"V_2=\frac{P_1 \times V_1}{P_2}= \frac{100 kPa \times 100 l}{800 kPa} = 12.5 l"